A certain printer cartridge has demonstrated a mean time usage of 9 hours and a standard deviation of 22 minutes. What time usage guarantee (in hours) should the manufacturer advertise in order to ensure that only 4% of the cartridges fail to meet the guaranteed time usage?
1
Expert's answer
2020-01-29T05:56:01-0500
"mean=9\cdot60\ min=540\ min""S.D.=22 min""\alpha/2=0.04, z_{\alpha/2}=1.7507"

"CI=mean\pm z_{\alpha/2}\cdot S.D"

"CI=540\pm 1.7507\cdot 22"

"CI=(501.5\ min; 578.5\ min )"

"CI=(8\ h \ 21.5\ min; 9\ h\ 38.5\ min )"